please help: (3x+2)^2=16
Answer by checkley77(12844)

(3x+2)^2=16
9x^2+12x+4=18
9x^2+12x+4-18=0
4x^2+12x-14=0
2(x^2+6x-7)=0
2(x+7)(x-1)=0
x+7=0
x=-7 ans.
x-1=0
x=1 ans.